A proton moving at 8.2×106 m/s enters a magnetic field of 0.38 Tesla, with the velocity vector at 90o to the magnetic field direction.

Calculate the magnitude of the force exerted on the particle.
Answer #1

given,

velocity of the proton = 8.2 * 10^6 m/s

magnetic field = 0.38 T

force = charge * velocity * magnetic field

force = 1.6 * 10^-19 * 8.2 * 10^6 * 0.38

force exerted on the particle = 4.986 * 10^-13 N